OpenStudy (anonymous):

The sum of two numbers is 51 . The larger number is 23 more than the smaller number. What are the numbers?

OpenStudy (lochana):

consider unknown numbers are 'x' and 'y'. and write two equations out of given information

OpenStudy (anonymous):

(23+x)+(23-y)=51?

OpenStudy (lochana):

no.. let's take first statement 'The sum of two numbers is 51' this means x + y = 51

OpenStudy (lochana):

second one is 'The larger number is 23 more than the smaller number' that means x is 23 more larger than y. x = y +23

OpenStudy (anonymous):

ok, so how do merge the 2 equations?

Directrix (directrix):

@splatter6 Let x be the larger number The smaller will be 50 - x The larger number is 23 more than the smaller number x = 50 - x + 23 x = 73 - x Solve for x. Subtract the value of x from 50 to get the second number. Expect fractions.